Circular No. 8315

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004bc IN NGC 3465 F. Manzini, Sozzago, Italy, reports his discovery of an apparent super-

nova (mag ∼ 16.5) on unfiltered CCD images taken with a 0.4-m telescope on Apr. 1.87–1.91 UT, noting the new object to be 3′′.5 west and 3′′.5 south of the center of NGC 3465, and adding that nothing is visible at this location on his image from Mar. 16 (limiting mag ∼ 19.0) or on Palomar Sky Survey images. Confirming unfiltered CCD images were obtained by Manzini, R. Crippa, and F. Fumagalli (0.4-m telescope) and by M. Amoretti (Sanremo, Italy, 0.16-m telescope) on Apr. 3.92, and by P. Holvorcem and M. Schwartz (0.81-m Tenagra telescope near Nogales, AZ) on Apr. 4.23. Holvocem reports the magnitude of SN 2004bc as 16.6 and provides the following precise position: α = 10^{h}59^{m}31^{s}.41, δ = +75^{o}11′28′′.9 (equinox 2000.0), which is 4′′.5 west and 4′′.3 south of the nucleus of NGC 3465.

SUPERNOVA 2004ac IN IC 4769 M. Phillips, M. Hamuy, and S. Gonzalez, Carnegie Observatories, re-

port that a spectrum (range 380–920 nm) of SN 2004ac (cf. IAUC 8295, 8296), obtained on Mar. 25.39 UT with the Las Campanas Observatory Dupont 2.5-m telescope (+ WFCCD), shows that this is a type-Ia supernova, ∼ 2.5 months past maximum light.

(2867) STEINS M. D. Hicks and J. M. Bauer, Jet Propulsion Laboratory, California

Institute of Technology; and and A. T. Tokunaga, Institute for Astronomy, University of Hawaii, write that they obtained BV RI photometry of minor planet (2867), a Rosetta flyby target, during Mar. 27 and 28 UT at the Table Mountain Observatory 0.6-m telescope. The minor planet exhibited a double-peaked lightcurve with a period of 6.06 ± 0.05 hr and amplitude 0.2 mag, corresponding to a maximum a/b triaxial ratio of 0.8. Assuming G = 0.15 for the phase-function slope parameter, averaged absolute magnitudes of 12.67 ± 0.02 in R and 13.18 ± 0.04 in V are derived. Hicks et al. add that the measured colors of B − R = +1.311 ± 0.026, V − R = +0.513 ± 0.029, and R − I = +0.440 ± 0.030 cannot rule out a D-type taxonomic classification but are more suggestive of an S-type asteroid.

2004 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green